Describe how you would determine an equation for a line given the ordered pairs for two points on that line.
Answer by mananth(16949)


Let the points be (x1,y1) & (x2,y2)
x1 y1 x2 y2
3 2 2 1

slope m =(y2-y1)/(x2-x1)
( 1 - 2 )/( 2 - 3 )
( -1 / -1 )
m= 1

Plug value of the slope and point ( 3 ,2 ) in
Y = m x + b
2.00 = 3 + b
b= 2 - 3
b= -1
So the equation will be
Y = 1 x -1
m.ananth@hotmail.ca